OpenStudy (anonymous):

Expand the brackets and simplify 2(x + 3) + 5(x-3)

OpenStudy (diyadiya):

2(x+3) + 5(x-3) first distribute 2 in 2(x+3) 2*x + 2*3 =2x + 6 can you do for 5(x-3) ?

OpenStudy (anonymous):

What?????????????????/

OpenStudy (diyadiya):

\[2(x+3) = 2*x + 2*3 = 2x + 6\] did you understand this?

OpenStudy (anonymous):

Yes

OpenStudy (diyadiya):

5(x-3) = 5*x - 5*3

OpenStudy (anonymous):

I kind of understand what you are trying to say but I'm a first timer at this

OpenStudy (anonymous):

Are you a teacher

OpenStudy (anonymous):

How long will it take for you to reply

OpenStudy (diyadiya):

ok a(b+c) = ab+ ac a(b-c) = ab - ac same way 2(x+3) here we distribute 2 2*x + 2*3 = 2x + 6 5(x-3) here we distribute 5 5*x - 5*3 5x - 15

OpenStudy (anonymous):

I understand now so after that what do you do

OpenStudy (diyadiya):

2(x + 3) + 5(x-3) = 2x + 6 + 5x -15 now we combine like terms we can only add numbers with same variables.

OpenStudy (anonymous):

So you add the x with the x and the numbers with the numbers

OpenStudy (diyadiya):

yes 2x+5x = 7x 6- 15= ?

OpenStudy (anonymous):

-9

OpenStudy (diyadiya):

yes so 2x + 6 + 5x -15 = 7x -9

OpenStudy (anonymous):

Thankyou Diyadiya if thats your real name

OpenStudy (diyadiya):

you're Welcome :) and no :D